Describe the translations applied to the graph of y 5 x2 to obtain a K graph of each quadratic function. Sketch the graph.

y=(x-5)^2+3
1
Expert's answer
2020-12-03T13:49:12-0500

First

The general equation

"y=a(x-h)\u00b2+k"

Where (h,k) is the turning point.

Y1=5(x)²

Y2=(x-5)²+3

Translation are


Horizontal shift by -5 units

Vertical shift by +3 units

Vertical compression by 1/5